Ms. Shruti is 45 years old and plans to retire at 60. Her life expectancy is 80 years. Ms.
Suman her Financial Planner, estimates that her client will require Rs.540,000 in the first
year after retirement. Inflation rate is 6% p.a. and the rate of return is 10% p.a. What will
be the savings per year required in order to meet this?
1
Expert's answer
2020-12-04T07:22:28-0500

this is an annuity

in the first case::

"540 000=\frac{FV\times0.10}{(1+0.10)^{20}-1)}"

FV=30 928 499.7

in the second case:

inflation adjusted rate:

"(\frac{1.1}{1.06}-1)\times100=6.796117"


"540 000=\frac{FV\times0.06796117}{(1+0.06796117)^{20}-1)}"

FV=21 650 909.3

30 928 499.7-21 650 909.3=9 277 590.36